a chemist wants to obtain 80 liters of a solution that is 12% hydrochloric acid. how many liters of 10% and 20% hydrochloric acid solutions should he mix to do so?
Answer by stanbon(75887)

a chemist wants to obtain 80 liters of a solution that is 12% hydrochloric acid. how many liters of 10% and 20% hydrochloric acid solutions should he mix to do so?
------------
Equation:
acid + acid = acid
0.10x + 0.20(80-x) = 0.12*80
--------
10x + 20*80 - 20x = 12*80
-10x = -8*80
x = 64 liters (amt. of 10% solution needed)
----
80-64 = 16 liters (amt. of 20% solution needed)
=========================
Cheers,
Stan H.